please help me....I think I am half way right.
79. If a vendor charges p dollars each for rugby shirts, then he expects to sell 2000-100p shirts at a tournament.
a.) find a polynomial R(p) that represents the total revenue when the shirts are p dollars each.
• Revenue = Price * Quantity
q = 2000-100p
R/P = 2000-100p
R = 2000p-100p^2
2000-200p = 0
200p=2000
p=10
is this part right??
b.) Find R(5),R(10) and R(20) <-----Need some help with this one

Answer by stanbon(75887)

If a vendor charges p dollars each for rugby shirts, then he expects to sell 2000-100p shirts at a tournament.
a.) find a polynomial R(p) that represents the total revenue when the shirts are p dollars each.
• Revenue = Price * Quantity
quantity = 2000-100p
price = p
---
Revenue = 2000p-100p^2
b.)
Find R(5) = 2000*5 - 100*5^2 = -15,000
------
R(10) = 2000*10 - 100*10^2 = 20,000 - 1,000,000 = -$800,000
------
R(20) = 2000*20 - 100*20^2 = 40,000 - 40,000 = 0
===================================================
Cheers,
Stan H.
==============